OpenStudy (anonymous):

Evaluate -(a + b), given values a = -4, b = - 12.

sam (.sam.):

-(a + b), given values a = -4, b = - 12 -(-4+-12) =16

OpenStudy (anonymous):

Sammi, do you know how to add and subtract negative numbers?

sam (.sam.):

-(-4+-12) 4+12=16 yes

OpenStudy (anonymous):

Lol, no I was talking to sammibaby

OpenStudy (anonymous):

b/c she needs to know that basic stuff before doing the problems she's doing right now.